# Learned enclosure method for experimental EIT data

### v1

@@ Abstract
Electrical impedance tomography (EIT) is a non-invasive imaging method with diverse applications, including medical imaging and non-destructive testing. The inverse problem of reconstructing internal electrical conductivity from boundary measurements is nonlinear and highly ill-posed, making it difficult to solve accurately. In recent years, there has been growing interest in combining analytical methods with machine learning to solve inverse problems. In this paper, we propose a method for estimating the convex hull of inclusions from boundary measurements by combining the enclosure method proposed by Ikehata with neural networks. We demonstrate its performance using experimental data. Compared to the classical enclosure method with least squares fitting, the learned convex hull achieves superior performance on both simulated and experimental data.
@@ 1 Introduction
This paper investigates the reconstruction method of combining the enclosure method and neural networks in EIT for experimental measurement data. We focus on reconstructing the convex hull of inclusions embedded in a constant background conductivity from boundary measurements. The previous work in [23] considered the continuum model in EIT and simulated data. To deal with experimental data, we need to consider the complete electrode model [6]. To show the validity of the proposed method, we demonstrate it with two-dimensional simulated data and experimental measurements from the open dataset published in [11]. The proposed neural network is able to recover the convex hull from noisy indicator function values significantly more accurately than the least squares approach used for comparison.
@@ 2 Enclosure method for EIT
$\Omega$$D$$D$$D$$x_{1}$$x_{2}$$\omega$
$x\cdot\omega=h_{D}(\omega)$
Figure 1: Left: support function of inclusion $D\subset\mathbb{R}^{2}$ with direction $\omega\in S^{1}$ by ${h_{D}(\omega)}:=\sup_{x\in D}x\cdot\omega.$ Here $\Omega$ is the unit disc. Right: real-world phantom with cylindrical plastic shapes immersed in saline (salt water). There are 16 electrodes in this EIT system.
@@
There are some assumptions on the function $h:\Omega\rightarrow\mathbb{R}$ ensuring that the conductivity $\sigma$ has a jump on the boundary $\partial D$. We assume that $D$ is a union of finitely many components with piecewise smooth boundaries.
@@ 2.1 Least squares fit
in the least squares sense yields the so-called least squares (LS) fit. For small $\tau$ values, the slope of the LS fit approximates the support function $h_{D}(\omega)$ in the direction $\omega$.
@@
Set $M=45$ and $N=10$, meaning that the indicator function is computed over 45 directions uniformly distributed in the interval $[0,2\pi]$ and over ten $\tau$ values ranging from $0.5$ to $5$ with step size $0.5$. Fixing $i=4$ allows us to compute $\frac{1}{2}\log|I_{\omega_{4}}(\tau_{j})|$ for ten $\tau_{j}$ values and fit a line through the points. The slope of this line approximates $h_{D}(\omega_{4})$. For an illustration see Figure 2.
@@
Figure 2: An example of the slope of the least squares fit approximating the true support function of the phantom on the right in the direction $\omega_{4}$. The values of $\frac{1}{2}\text{log}|I_{\omega_{4}}(\tau_{j})|$, $j=1,\dots,10$, are shown on the left as dots. The LS fit is plotted as a dotted and $h_{D}(\omega_{4})\tau+b$ as a solid line. The true support function $h_{D}(\omega_{4})$ corresponds to the distance of the probing half-plane from the origin, shown on the right in dark gray. The slope of the LS fit approximates this distance but overestimates it, as can be seen from the light gray half-plane. To get the full convex hull of inclusions, this process is repeated for all directions $\omega_{i}$.
@@ 2.2 Continuum measurement model
$I_{\omega}(\tau)=\int_{\partial\Omega}f_{\omega}(\,\cdot\,;\tau)\Big{(}\Lambda_{\sigma}-\Lambda_{1})\overline{f_{\omega}(\,\cdot\,;\tau)}\Big{)}dS,$ (9)
@@ 3.1 Training data
We simulated 10 000 training data phantoms consisting of the unit disk with 1 to 3 disjoint elliptical inclusions. The distance between inclusions is at least 0.05 and the distance between each inclusion and the unit circle is at least 0.1, but otherwise the placement of the inclusions is random. Each elliptical inclusion has a random orientation and the length of the semi-major axis $R$ is uniformly distributed in the interval $[0.05,0.3]$, and the length of the semi-minor axis is uniformly distributed in the interval $[0.05,R]$. The conductivity values of the inclusions were chosen to be either more conductive or resistive than the constant background conductivity value 1, with probability $1/2$. The specific conductive and resistive values are uniformly distributed in the interval $[2,4]$ for conductive inclusions and $[0.2,0.85]$ for resistive inclusions.
After generating the conductivity phantoms, we simulated the DN map $\boldsymbol{\Lambda}_{\sigma}$ for each phantom by solving the forward EIT problem as explained in [21, Section 13.2] using the continuum model. We added random noise of magnitude $10^{-4.5}$ to the DN maps and used equation (9) to compute the indicator function $I_{\omega_{i}}(\tau_{j})$ for $\omega_{i}$ and $\tau_{j}$ defined as in the example 5. Although $\tau$ is taken to infinity in equation (3) to reveal the support function, severely restricting $\tau$ is necessary in practical settings because measurement noise makes the indicator function unstable with higher $\tau$ values [23].
In accordance with the LS fit, the values of the indicator functions were used to compute the points
@@
for each $\omega_{i}$ and $\tau_{j}$. The resulting values were gathered into a $10\times 45$ matrix with rows corresponding to the $\tau$ values and columns corresponding to the 45 directions of $\omega$. To accentuate the periodic nature of the indicator function, we followed a similar padding approach as in [23], so that each indicator function matrix is reflect-padded to size $30\times 50$.
@@ 3.2 Network architecture
We use a shallow feed-forward convolutional network implemented with MATLAB’s Deep Learning Toolbox to learn the support vectors based on the inputs. See Figure 4 for an illustration of the network architecture. In the convolutional layers, we use a three-pixel zero padding. The network was trained using the half-mean-squared-error loss function for 200 epochs. The initial learning rate was $0.01$, which was reduced every 5 epochs by a factor of $0.2$.
@@ 4 Results
We use the above neural network model to learn the support vectors based on the inputs, and compute the learned convex hull of the inclusions from each learned support vector.
The true support function, computed directly from the conductivity phantom, can be used to form a fairly accurate approximation of the true convex hull. We use this approximated hull as a ground truth. In a similar manner, the LS fit (4) can be used to compute the least squares convex hull, a rough approximation of the true convex hull. In order to assess the performance of the neural network, we compare the learned convex hulls, least squares convex hulls and ground truths. We test the proposed method of learned convex hulls with both simulated and experimental data. In the case of simulated data, we compare the errors of the least squares convex hulls and learned convex hulls relative to the ground truths. We have no access to the ground truths of the experimental data, so we simply compare the learned convex hulls with the least squares convex hulls visually.
@@ 4.1 Simulated data
Histogram comparisons of the relative errors of the least squares solution versus the proposed method are presented in Figure 5, and Figure 4.1 shows examples of the simulated data results. The example cases were chosen to demonstrate the performance of the method in a diverse set of conductivity phantoms.
(a) Convex hull relative errors (%)
(b) False positives (%)
(c) False negatives (%)
Figure 5: Histograms depicting the relative errors of the learned convex hulls (blue) and least squares convex hulls (red) over the simulated testing data set.
Case Least squares Learned Ground truth Support vectors
@@
Figure 0: Comparison of the convex hulls and support vectors of the simulated phantoms, computed using LS (dotted line), learned approach (dashed line), and the ground truth (solid line). The error relative to the ground truth is shown below each phantom.
@@ 4.2 Experimental data
We use a total of 22 experimental phantoms from the open dataset [11] to test the method. The experimental setup consists of a circular cylinder tank (diameter 28cm) filled with saline, with either plastic (resistive) or metal (conductive) inclusions. Each phantom has 1-3 circular, triangular, or square inclusions. The KIT4 measurement system [20] at the University of Eastern Finland used to collect the measurements has 16 electrodes. For this study, we have omitted the cases from the dataset with the annular foam layer on the boundary and the pumpkin. See Figures -11 and -20 for the experimental data results. The results follow the numbering scheme in the original open data set and they are presented roughly in ascending order of difficulty.
Case Phantom Hull Case Phantom Hull
@@
Figure -11: Learned convex hull (dark gray) versus the least squares convex hull (light gray) for the experimental phantoms.
Case Phantom Hull Case Phantom Hull
@@
Figure -20: Learned convex hull (dark gray) versus the least squares convex hull (light gray) for the experimental phantoms.
@@ 5 Discussion
5.1 Performance on simulated data.
For simulated data, the learned convex hull surpasses the least squares hull in accuracy for all 1000 conductivity phantoms in the test set. Based on the error analysis in the histograms, using the learned convex hull significantly decreases the relative errors within the simulated testing set (Figure 5).
As can be seen in the histogram of false positives (Figure 5(b)), the least squares approach is highly prone to overestimation of the hull. This can already be seen in Figure 2 with a noise-free example, and adding noise to the DN maps further exacerbates the issue. On the other hand, the learned hull captures the inclusions quite well, but sometimes underestimates the size of the hull, leaving some parts of the inclusions outside, as can be seen in Figure 4.1 in the examples 4.1 and 4.1 where the jump condition fails, and interestingly also in example 4.1. However, based on the histogram of false negatives (Figure 5(c)), this problem is also present and more serious with the LS approach for some cases in the test set.
@@ 5.2 Performance on experimental data
Even though the network was trained with simulated continuum model data, it generalizes well for experimental data, as can be seen in the results in Figures -11 and -20. The figures show that the learned convex hull produces a more accurate reconstruction of the convex hull compared to the least squares convex hull for all phantoms, except for case -20 where the learned hull predicts that there is no inclusion. Since the least squares hull is also having trouble with this case, the cause of the inaccuracy is most likely already present in the indicator function, and thus in the input given to the neural network. Resistive inclusions are generally more difficult to detect, which can be a factor in the poor results for both approaches.
Due to distinguishability issues [18], inclusions further away from the boundary are exponentially more difficult to detect from the DN map, which might contribute to the lackluster results in cases -20 and -20. The inclusions are also very small, which adds to the difficulty.
For phantoms with both plastic and metal inclusions where the jump condition fails, the learned convex hull is able to find the location of the inclusions quite well, except in case -20 where the conductive and resistive inclusions are very close to each other. These cases are not covered by the theory and the LS hull also produces poor results, especially for the aforementioned case. Thus, it is quite remarkable to even have coarse estimates for these cases with experimental data.
@@ 5.3 Influence of network size and training data
When training the neural network, we tested different architectures and training set sizes. Having more convolutional layers improved the results on simulated data, but made the network overfit. When testing with experimental data, the larger networks made the convex hulls too small or even completely disappear. Increasing the amount of training data did not improve the results, likely because the number of trainable parameters is quite small. This is expected, since we don’t need high expressivity in the network as the learned task is fairly low-dimensional.

### v3

@@ Abstract
Electrical impedance tomography (EIT) is a non-invasive imaging method with diverse applications, including medical imaging and non-destructive testing. The inverse problem of reconstructing internal electrical conductivity from boundary measurements is nonlinear and highly ill-posed, making it difficult to solve accurately. In recent years, there has been growing interest in combining analytical methods with machine learning to solve inverse problems. In this paper, we propose a method for estimating the convex hull of inclusions from boundary measurements by combining the enclosure method proposed by Ikehata with neural networks. We demonstrate its performance using experimental data. Compared to the classical enclosure method with least squares fitting, the learned hull achieves superior performance on both simulated and experimental data.
@@ 1 Introduction
This paper investigates the reconstruction method of combining the enclosure method and neural networks in EIT for experimental measurement data. We focus on reconstructing the convex hull of inclusions embedded in a constant background conductivity from boundary measurements. The previous work in [25] considered the continuum model in EIT and simulated data. To show the validity of the proposed method, we demonstrate it with two-dimensional simulated data and experimental measurements from the open dataset published in [12]. The proposed neural network is able to recover the convex hull from noisy indicator function values significantly more accurately than the least squares approach used for comparison. Potential real-world applications for the proposed method could include skin cancer detection [23], and detecting cracks in the interior of a conductive object or in conductive paint on the surface of a structure [11, 9].
@@ 2 Enclosure method for EIT
$\Omega$$D$$D$$D$$x_{1}$$x_{2}$$\omega$Refer to caption
Figure 1: The image on the right shows an experimental phantom with plastic shapes placed in saline (salt water). On the left is a simulated phantom resembling the experimental setting. The image shows the line determined by the support function ${{h_{D}(\omega)}:=\sup_{x\in D}x\cdot\omega}$ in the direction $\omega\in S^{1}$ for the inclusions $D\subset\Omega$, where $\Omega$ is the unit disk.
@@
We assume that $D$ is a union of finitely many components with piecewise smooth boundaries. In [16], it is assumed that $h:\Omega\rightarrow\mathbb{R}$ satisfies the so-called jump condition along with other assumptions, ensuring that the conductivity $\sigma$ has a jump on the boundary $\partial D$. In practice, the jump condition means that the components of $D$ should be either all conductive or all resistive compared to the background conductivity. The theory does not cover cases where $D$ consists of both conductive and resistive inclusions, and using the enclosure method in such cases can lead to errors in the reconstructed convex hull [25].
@@ 2.1 Least squares fit
in the least squares sense yields the so-called least squares (LS) fit. For small $\tau$ values, the slope of the least squares fit approximates the support function $h_{D}(\omega)$ in the direction $\omega$.
@@
Figure 2: An example of the slope of the least squares fit approximating the true support function of the phantom on the right in the direction $\omega_{4}$. The values of $\frac{1}{2}\text{log}|I_{\omega_{4}}(\tau_{j})|$, $j=1,\dots,10$, are shown on the left as dots. The LS fit is plotted as a dotted and $h_{D}(\omega_{4})\tau+b$ as a solid line. The true support function $h_{D}(\omega_{4})$ corresponds to the distance of the probing half-plane from the origin, shown on the right in dark gray. The slope of the least squares fit approximates this distance but overestimates it, as can be seen from the light gray half-plane. To get the full convex hull of inclusions, this process is repeated for all directions $\omega_{i}$.
Set $M=45$ and $N=10$, meaning that the indicator function is computed over 45 directions uniformly distributed in the interval $[0,2\pi]$ and over ten $\tau$ values ranging from $0.5$ to $5$ with step size $0.5$. As an example, fixing $i=4$ allows us to compute $\frac{1}{2}\log|I_{\omega_{4}}(\tau_{j})|$ for ten $\tau_{j}$ values and fit a line through the points. The slope of this line approximates $h_{D}(\omega_{4})$. See Figure 2 for an illustration.
@@ 2.2 Continuum measurement model
[…]
@@ 3.1 Training data
We simulated 10 000 training data phantoms consisting of the unit disk with 1 to 3 disjoint elliptical inclusions. The distance between inclusions is at least 0.05 and the distance between each inclusion and the unit circle is at least 0.1, but otherwise the placement of the inclusions is random. Each elliptical inclusion has a random orientation and the length of the semi-major axis $R$ is uniformly distributed in the interval $[0.05,0.3]$, and the length of the semi-minor axis is uniformly distributed in the interval $[0.05,R]$. The conductivity values of the inclusions were chosen to be either more conductive or resistive than the constant background conductivity value 1, with probability $1/2$. The specific conductive and resistive values are uniformly distributed in the interval $[2,4]$ for conductive inclusions and $[0.2,0.8]$ for resistive inclusions. Notably, the training data set contains both phantoms that satisfy the jump condition and phantoms that don’t.
After generating the conductivity phantoms, we simulated the DN map $\boldsymbol{\Lambda}_{\sigma}$ for each phantom and the unit disk without inclusions by solving the forward EIT problem as explained in [22, Section 13.2], using the continuum model and finite element method in 2D. We used a finite element mesh with 33025 points and 65536 triangles with the maximum edge length 0.0123, providing solutions with 6 significant digits of accuracy. We added white Gaussian noise of magnitude $10^{-4.5}$ to the resulting Neumann-to-Dirichlet (ND) matrices. The magnitude of the noise was chosen so that significant information in the ND matrices is not lost. We then inverted the matrices to obtain approximations of the DN maps and used equation (9) to compute the indicator functions $I_{\omega_{i}}(\tau_{j})$ for $\omega_{i}$ and $\tau_{j}$ defined as in the example (5). Although $\tau$ is taken to infinity in equation (3) to reveal the support function, severely restricting $\tau$ is necessary in practical settings because measurement noise makes the indicator function unstable with higher $\tau$ values [25].
In accordance with the least squares fit, the values of the indicator functions were used to compute the points
@@
for each $\omega_{i}$ and $\tau_{j}$. The resulting values were gathered into a $10\times 45$ matrix with rows corresponding to the $\tau$ values and columns corresponding to the 45 directions $\omega$. To accentuate the periodic nature of the indicator function, we followed a similar padding approach as in [25], so that each indicator function matrix is reflect-padded to size $30\times 50$.
@@ 3.2 Network architecture
We use a shallow feed-forward convolutional network implemented with MATLAB’s Deep Learning Toolbox to learn the support vectors based on the inputs. See Figure 4 for an illustration of the network architecture. In the convolutional layers, we use a three-pixel zero padding. The network was trained using the half-mean-squared-error loss function for 200 epochs. The optimization method was Stochastic gradient descent with momentum of $0.9$, with an initial learning rate of $0.01$, which was reduced every 5 epochs by a factor of $0.2$.
@@ 4 Results
We test the proposed method of learned hulls with both simulated and experimental data. We use the above neural network model to learn the support vectors based on the inputs, and compute the learned hull of the inclusions from each learned support vector.
For experimental data, we created phantoms by segmenting the objects in the photographs of the measurements available in [12]. We computed the true support functions for both simulated and experimental data directly from the simulated or segmented conductivity phantoms, and used the true support functions to compute the true convex hulls. We use the true convex hulls as a ground truths.
As demonstrated in Figure 2, the least squares fit (4) can be used to compute the least squares hull, a rough approximation of the true convex hull. In order to assess the performance of the neural network, we compare the errors of the least squares hulls and learned hulls relative to the ground truths.
Denote the true convex hull by $\mathcal{C}\subset\mathbb{R}^{2}$ and the reconstructed convex hull by $\mathcal{B}\subset\mathbb{R}^{2}$. To measure the quality of $\mathcal{B}$ as an approximation to $\mathcal{C}$ quantitatively, we use the relative error formula
$\frac{|\mathcal{C}\setminus\mathcal{B}|+|\mathcal{B}\setminus\mathcal{C}|}{|\Omega|}\cdot 100\%,$ (22)
[…]
Refer to $\mathcal{B}$$\mathcal{C}$$\mathcal{C}\setminus\mathcal{B}$(false negatives)$\mathcal{B}\setminus\mathcal{C}$(false positives)
Figure 5: We measure the quality of reconstructions of convex hulls using formula (22). This picture illustrates how the error can be seen as a sum of false negatives and false positives.
@@ 4.1 Simulated data
Histogram comparisons of the relative errors of the least squares solution versus the proposed method are presented in Figure 6, and Figure 1 shows examples of the simulated data results. The example cases were chosen to demonstrate the performance of the method in a diverse set of conductivity phantoms.
(a) Convex hull relative errors
(b) False positives
(c) False negatives
Figure 6: Histograms showing the distribution of the relative errors of the learned hulls (blue) and least squares hulls (red) in the simulated testing data set. The horizontal axis represents the relative error (22) and the vertical axis represents the frequency. Most of the learned hulls have a relative error that falls under 10%, whereas the relative errors of the least squares hulls range from 7% to 78%. False positives account for most of the errors, meaning that both approaches tend to overestimate the hull.
Case Support vectors Ground truth Learned Least squares
I Refer to caption Refer to caption
@@ VI Refer to caption Refer to caption
Figure 1: Comparison of the support vectors and the hulls of the simulated phantoms, computed using LS (dotted line), learned approach (dashed line), and the ground truth (solid line). The error relative to the ground truth is shown below each phantom.
@@ 4.2 Experimental data
We use a total of 22 experimental phantoms from the open dataset [12] to test the method. The experimental setup consists of a circular cylinder tank (diameter 28cm) filled with saline, with either plastic (resistive) or metal (conductive) objects. Each phantom has 1-3 circular, triangular, or square objects. The KIT4 measurement system [21] at the University of Eastern Finland used to collect the measurements has 16 electrodes.
We use conductivity phantoms created from the photographs of the measurement setting as ground truths. To combat the distortion in the images caused by saline and the camera lens, we obtained ground truths by segmenting the objects in the images according to their highest point. Since the support function detects jumps on the boundaries of the inclusions and not the actual conductivities, it is not necessary for the inclusions to match the true conductivities of the plastic and metal objects. With this in mind, we assigned the conductive inclusions a value of 10 and the resistive inclusions a value of 0.1. The conductivity of the background was set to 1. We created a conductivity phantom based on the segmentation, and computed true support functions and true convex hulls according to (2). See Figure -16 for an example of the segmentation.
See Figures -4--17 for the experimental data results. The results follow the numbering scheme from the original open data set and they are presented roughly in ascending order of difficulty.
Case Phantom Ground truth Learned Least squares
2.1 Refer to caption
@@ 2.6 Refer to caption
92.5%
Figure -4: Comparison of the ground truth, learned and least squares hulls of the experimental phantoms. The error relative to the ground truth is shown below each phantom.
Case Phantom Ground truth Learned Least squares
3.1 Refer to caption
@@ 3.5 Refer to caption
64.4%
3.6 Refer to caption
[…]
69.6%
Figure -9: Comparison of the ground truth, learned and least squares hulls of the experimental phantoms. The error relative to the ground truth is shown below each phantom.
Case Phantom Ground truth Learned Least squares
4.1 Refer to caption
[…]
69.6%
4.2 Refer to caption
[…]
80.5%
4.3 Refer to caption
[…]
90.0%
4.4 Refer to caption
[…]
117.4%
5.1 Refer to caption
[…]
21.7%
5.2 Refer to caption
[…]
64.6%
Figure -14: Comparison of the ground truth, learned and least squares hulls of the experimental phantoms. The error relative to the ground truth is shown below each phantom.
Case Phantom Ground truth Learned Least squares
1.1 Refer to caption
[…]
92.6%
1.2 Refer to caption
[…]
62.6%
1.3 Refer to caption
[…]
73.2%
1.4 Refer to caption
[…]
91.1%
Figure -17: Comparison of the ground truth, learned and least squares hulls of the experimental phantoms. The error relative to the ground truth is shown below each phantom.
@@ 5 Discussion
5.1 Performance on simulated data
For simulated data, the learned hull surpasses the least squares hull in accuracy for all 1000 conductivity phantoms in the test set. Based on the error analysis in the histograms, using the learned hull significantly decreases the relative errors within the simulated testing set (Figure 6), even though the testing set contains cases that violate the jump condition and are therefore not covered by theory.
As can be seen in the histogram of false positives (Figure 6(b)), the least squares approach is highly prone to overestimation of the hull. This can already be seen in Figure 2 with a noise-free example, and adding noise to the DN maps further exacerbates the issue. On the other hand, the learned hull captures the inclusions quite well, but sometimes underestimates the size of the hull, leaving some parts of the inclusions outside, as can be seen in Figure 1 in the examples 1 and 1 where the jump condition fails, and interestingly also in example 1. However, based on the histogram of false negatives (Figure 6(c)), this problem is also present and more serious with the least squares approach for some cases in the test set.
@@ 5.2 Performance on experimental data
Even though the network was trained with simulated continuum model data, it generalizes well for experimental data, as can be seen in the results in Figures -4, -9, -14, and -17. The figures show that the learned hull is a more accurate reconstruction of the ground truth compared to the least squares hull for all phantoms except for the case -9, where both reconstructions underestimate the hull. It could also be argued that the least squares approach performs better in the case -17, because the learned hull predicts that there is no inclusion at all. Since the least squares hull is also having trouble with the case -17, the cause of the inaccuracy is most likely already present in the indicator function, and thus in the input given to the neural network. Resistive inclusions are generally more difficult to detect, which can be a factor in the poor results for both approaches.
Due to distinguishability issues [19], inclusions further away from the boundary are exponentially more difficult to detect from the DN map, which might contribute to the lackluster results in cases -17 and -17. The inclusions are also very small, which adds to the difficulty.
When there are multiple inclusions that satisfy the jump condition, it appears that both the learned and the least squares approach tend to underestimate the hull of resistive inclusions and overestimate the hull of conductive inclusions. This is best visualized in Figure -9. For phantoms with both plastic and metal inclusions where the jump condition fails, the learned hull is able to find the location of the inclusions quite well, except in case -14 where the conductive and resistive inclusions are very close to each other. These cases are not covered by the theory and the least squares hull also produces poor results, especially for the aforementioned case. Thus, it is quite remarkable to have even coarse estimates for these cases with experimental data.
Segmenting the objects according to their highest point makes the segmentation consistent across images, but it introduces errors into the ground truth since the objects have different heights. This is especially true for the cases -9, -14, -14, and -17 that feature a plastic cylinder that is significantly taller than the other segmented objects. As an example, a closer look at the segmentation of the case 3.3. is shown in Figure -16. However, due to the way the relative error is calculated (22), a small error in the location of one inclusion is likely to have a minimal effect on the relative error over the whole tank.
Figure -16: The image on the left shows an example of the segmentation and the image on the right shows the resulting convex hull that is used as ground truth.
@@ 5.3 Influence of network size and training data
When training the neural network, we tested different architectures and training set sizes. Having more convolutional layers improved the results on simulated data, but made the network overfit. When testing with experimental data, the larger networks made the learned hulls too small or even completely disappear. Increasing the amount of training data did not improve the results, likely because the number of trainable parameters is quite small. This is expected, since we don’t need high expressivity in the network as the learned task is fairly low-dimensional.